OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(-5, 1), C(0, -4), D(5, 1) and E (0, 6) by showing that all four sides are equal.

OpenStudy (shayaan_mustafa):

Hi drpick :) How are you dear?

OpenStudy (shayaan_mustafa):

It is very easy. Find distances from one point to another and you will find that all distances are equal.

OpenStudy (shayaan_mustafa):

Hence in Rhombus all sides are equal.

OpenStudy (shayaan_mustafa):

Find distances between BC CD DE EB

OpenStudy (shayaan_mustafa):

Use this formula to find result \[\Large \sqrt{(y_2-y_1)^2+(x_2-x_1)^2}\]

OpenStudy (anonymous):

\[ BC^2=(-4-1)^2+(0+5)^2=50\\ CD^2=(5+0)^2++(1+4)^2=50 \] Find the square of the other two sides in the same way.